alones (alones):

If the legs of a right triangle measure 5 and 10, what is the length of the hypotenuse?

alones (alones):

i'm not sure tho but is is 5√ 5

OpenStudy (shastri007):

Using Pyth Thrm, c^2 = a^2 + b^2 c^2 = (5)^2 + (10)^2 c^2 = 25 + 100 c = Squareroot of 125 Which 5 root 5 Indeed correct.! Good job

alones (alones):

thanks! ;)

OpenStudy (across):

Let\[f(x)=-\frac{1}{2}x+5.\]Then\[f'(x)=-\frac12\]and\[\int_0^{10}\sqrt{1+\left(-\frac12\right)^2}\,dx=\left.\frac{\sqrt5}{2}x\right|_0^{10}=5\sqrt5.\]

OpenStudy (shastri007):

Your welcome :)

OpenStudy (nubeer):

@across your way is a bit tough for her i suppose as she haven't done integration yet.. :3

OpenStudy (across):

I'm being silly. :-P I'm on a rampage of overkilling simple questions.

alones (alones):

LOOOL CROSS

OpenStudy (nubeer):

oh hahaha .. then go ahead kill some more :p